OpenStudy (anonymous):

Evaluate the function rule for the given value. f(x) = 10^2 for x = 2

OpenStudy (australopithecus):

are you sure this is how it is written?

OpenStudy (australopithecus):

you can assume that the function f(x) = 10^(2) where 10^(2) is the out put thus you have to devise a function that contains x that will give you 10^(2) thus you can assume f(x) = 10^(x) thus f(2) = 10^(2)

OpenStudy (australopithecus):

this is a weird question thought that is for sure